OpenStudy (anonymous):

The ages of two exercise classes at a local fitness center are listed below. Which statement about the data is true? Class A: 37, 36, 32, 41, 42, 36, 44, 45 Class B: 26, 40, 38, 36, 42, 40, 40, 36 a. The mean age is the same for both classes. b. The range of ages is greater in Class A that in Class B. c. The mode of the ages for both classes is 36. d. The median age is the same for both groups.

OpenStudy (texaschic101):

put the numbers in order... Class A -- 32, 36, 36, 37, 41,42, 44, 45 median is (37 + 41)/2 = 78/2 = 39 mode is 36 range is 45 - 32 = 13 mean is (32 + 36(2) + 37 + 41 + 42 + 44 + 45)/8 = 313/8 = 39.13 Class B --26, 36,36,38,40, 40, 40, 42 median is (38 + 40)/2 = 78/2 = 39 mode is 40 range is 42 - 26 = 16 mean is (26 + 36(2) + 38 + 40(3) + 42)/8 = 298.8 = 37.25 Its the median...they are both the same

OpenStudy (anonymous):

thank you!(:

OpenStudy (texaschic101):

sure thing...sorry it took so long

OpenStudy (anonymous):

its okay it helped me (: